OpenStudy (anonymous):

determine the tan of 810°

Parth (parthkohli):

\( \color{Black}{\Rightarrow 810^{\circ} \Longrightarrow 720^{\circ} + 90^{\circ}}\) 720 degrees is just 0 degrees. \( \color{Black}{\Rightarrow \tan 810 = \tan 90 }\)

OpenStudy (anonymous):

undefined

Parth (parthkohli):

\( \color{Black}{\Rightarrow \tan 90 = \infty}\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow \tan 90 = \large {\sin 90 \over \cos 90} = {1 \over 0} = \infty}\)

Parth (parthkohli):

@dorkkk Did you get it?

OpenStudy (anonymous):

why does it become tan90?

Parth (parthkohli):

See my first reply.

Parth (parthkohli):

810 degrees is just 2 revolutions and then 1/4 revolution right? 2 revolutions are just 0 degrees, and 1/4 revolution is 90.

OpenStudy (anonymous):

oh, okay i see that when i plug it in to my calculator is undefined :P

Parth (parthkohli):

Yep.. tan 90 is always undefined :)